Circular No. 7643

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cn IN IC 4758 R. Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag ∼ 14.8) on unfiltered CCD images taken on June 11.95 and 12.85 UT with a 0.30-m reflector. SN 2001cn is located at α = 18^{h}46^{m}17^{s}.84, δ = −65^{o}45′41′′.8 (equinox 2000.0), which is 2′′.6 west and 17′′.9 south of the nucleus of IC 4758. The new object does not appear on a Digital Sky Survey image (U.K. Schmidt 1.2-m telescope) taken on 1976 May 4.75 (limiting mag ∼ 21.0). R. Santallo, Faaa, Tahiti, French Polynesia, reports a confirmation of SN 2001cn at mag ∼ 14.8 on unfiltered CCD images taken on June 13.531 with a 0.2-m reflector, and he provides the following position end figures: 17^{s}.88, 41′′.8 (or 1′′.8 west and 17′′.4 of the galaxy nucleus).

SUPERNOVA 2001co IN NGC 5559 A. B. Aazami and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.5) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on June 11.3, 12.3, and 13.3 UT. The new object is located at α = 14^{h}19^{m}11^{s}.80, δ = +24^{o}47′42′′.9 (equinox 2000.0), which is 13′′.5 west and 12′′.6 south of the nucleus of NGC 5559. A KAIT image taken on May 24.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001cl IN NGC 7260 S. Benetti, G. Altavilla, E. Cappellaro, A. Pastorello, and M. Turatto,

Osservatorio Astronomico di Padova; and J. Pritchard, European Southern Observatory (ESO), write: “A fully reduced CCD spectrum (range 330– 920 nm, resolution 1.0 nm) of SN 2001cl (cf. IAUC 7641), obtained on June 12.32 UT with the Danish 1.54-m telescope (+ DFOSC) at ESO, is that of a type-II supernova, 3–4 weeks after maximum light. The spectrum shows a broad Hα emission with a shallow P-Cyg absorption having a minimum measured at 651 nm. Adopting the NASA/IPAC Extragalactic Database recession velocity for NGC 7260 of 4901 km/s, this indicates an expansion velocity of ∼ 6500 km/s. Other strong absorption features are measured at 578 (Na I D + He I); 518 and 503 (Fe II); and 485 nm (Hβ).”

2001 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green